Analyzing Transactions Using the Financial Statement Effects Template
Following are selected transactions of Mogg Company. Record the effects of each using the financial statement effects template.

  1. Shareholders contribute $10,000 cash to the business in exchange for common stock.
  2. Employees earn $500 in wages that have not been paid at period-end.
  3. Inventory of $3,000 is purchased on credit.
  4. The inventory purchased in transaction 3 is sold for $4,500 on credit.
  5. The company collected the $4,500 owed to it per transaction 4.
  6. Equipment is purchased for $5,000 cash.
  7. Depreciation of $1,000 is recorded on the equipment from transaction 6.
  8. The Supplies account had a $3,800 balance at the beginning of this period; a physical count at period-end shows that $800 of supplies are still available. No supplies were purchased during this period.
  9. The company paid $12,000 cash toward the principal on a note payable; also, $500 cash is paid to cover this note's interest expense for the period.
  10. The company receives $8,000 cash in advance for services to be delivered next period.

Use negative signs with your answers, when appropriate.
Hint: For transaction 4, enter the net effect amount for balance sheet answers.

Balance Sheet

Income Statement

Transaction Cash
Asset
+ Noncash
Assets
= Liabilities + Contributed
Capital
+ Earned
Capital

Revenue

-

Expenses

=

Net Income
1 = - =
2 = - =
3 = - =
4 = - =
5 = - =
6 = - =
7 = - =
8 = - =
9 = - =
10 = - =
Answer #1
Balance Sheet Income Statement
Transaction Cash + Noncash = Liabilities + Contributed + Earned Revenue - Expenses = Net Income
Asset Assets Capital Capital
1 $ 10,000.00 + $            -   = $               -   + $10,000.00 + 0 0 - 0 = 0
2 $               -   + $            -   = $       500.00 + 0 + $   (500.00) 0 - $   500.00 = $   (500.00)
3 $               -   + $ 3,000.00 = $    3,000.00 + 0 + 0 0 - 0 = 0
4 $               -   + $ 1,500.00 = $               -   + 0 + $ 1,500.00 $4,500.00 - $3,000.00 = $ 1,500.00
5 $    4,500.00 + $(4,500.00) = $               -   + 0 + 0 0 - 0 = 0
6 $ (5,000.00) + $ 5,000.00 = $               -   + 0 + 0 0 - 0 = 0
7 $               -   + $(1,000.00) = $               -   + 0 + $(1,000.00) 0 - $1,000.00 = $(1,000.00)
8 $               -   + $(3,000.00) = $               -   + 0 + $(3,000.00) 0 - $3,000.00 = $(3,000.00)
9 $(12,500.00) + $            -   = $(12,000.00) + 0 + $   (500.00) 0 - $   500.00 = $   (500.00)
10 $    8,000.00 + $            -   = $    8,000.00 + 0 + 0 0 - 0 = 0